OpenStudy (anonymous):

s^2-9/s^2-6s+9 simplify by removing factors of 1

OpenStudy (anonymous):

=(s+3)(s-3)/(s-3)^2 =(s+3)/(s-3)

OpenStudy (anonymous):

wants in fractionized form

jhonyy9 (jhonyy9):

us (a*2 -b*2)/(a-b)*2 =((a-b)(a+b))/(a-b)*2 after you have simplified with (a-b) you get (a+b)/(a-b)